# Fibred surfaces and their unitary rank

### v1

@@ 2 Fibred surfaces and their geography
For arbitrary $b$ he proved the bound
@@
He also conjectured that the bound (2.18) holds for any non-trivial fibration. Serrano [49] proved that if $f$ is isotrivial (but not trivial), then (2.18) holds. Cai in [15] proved that if $f$ is non-isotrivial and the general fibre is either hyperelliptic or bielliptic, the same bound holds. In the article that inspired our present work Konno improved the bound (2.18) to:
$q_{f}\leq g\frac{5g-2}{6g-3}$ (2.20)
Pirola gave a counterexample of genus $3$ to Xiao’s conjecture in [46], by constructing for non-isotrivial fibrations a higher Abel-Jacobi map. More counterexamples have been found later by Albano and Pirola in [1].
In [6] Barja, González and Naranjo proved the following: if $f$ is non-isotrivial, then $q_{f}\leq g-c_{f}$; moreover, they made the following modified conjecture:
@@
Favale-Naranjo-Pirola [24] have proven the stronger inequality $q_{f}\leq g-c_{f}-1$ for families of plane curves of degree $d\geq 5$.
It is natural to ask wether these bounds can be extended to the unitary rank $u_{f}$.
In the paper [30] the author with González and Torelli proved the inequality $u_{f}\leq g-c_{f}.$ Moreover, if the general fibre is a plane curve of degree $d\geq 5$, then $u_{f}\leq g-c_{f}-1.$
@@
If $\mathcal{U}$ has finite monodromy, then up to a base change it becomes trivial, and so coincides with $q_{f}$. However, the unitary summand $\mathcal{U}$ of the Hodge bundle can have infinite monodromy as proved by Catanese and Dettweiler in [16]. Moreover, modified Xiao’s conjecture does not hold for $u_{f}$: this follows from [16, Remark 38 ] and from the construction of Lu [41]. See also [30, Conjecture 1.1].
@@ 3 Main technical results
3.1 The Harder-Narashiman sequence of a vector bundle
Let us briefly recall the notion of the Harder-Narasimhan sequence of a vector bundle $\mathcal{E}$ over a curve $B$ (see [31]). It is the unique filtration of subbundles
$0=\mathcal{E}_{0}\subset\mathcal{E}_{1}\subset\ldots\subset\mathcal{E}_{l}=\mathcal{E}$
@@ 3.1 The Harder-Narashiman sequence of a vector bundle
  • for any $i=0,\ldots l$ $\mathcal{E}_{i}/\mathcal{E}_{i-1}$ is a $\mu$-semistable vector bundle;
  • if we set $\mu_{i}:=\mu(\mathcal{E}_{i}/\mathcal{E}_{i-1})$, we have that $\mu_{i}>\mu_{i-1}$.
Note that $\mu_{1}>\mu(\mathcal{E})>\mu_{l}$, unless $\mathcal{E}$ is $\mu$-semistable, in which case $1=l$ and $\mu(\mathcal{E})=\mu_{1}$. Note moreover that, setting $r_{0}:=0$, we can express the degree of $\mathcal{E}$ as a combinations of the $\mu_{i}$’s and $r_{i}$’s:
@@
Let us call $H$, respectively $\Sigma$ a tautological divisor, respectively a fibre on the projective bundle $\pi\colon\mathbb{P}_{B}(\mathcal{E})\to B$. Recall that $\mbox{{Pic}}(\mathbb{P}_{B}(\mathcal{E}))=\pi^{*}\mbox{{Pic}}(B)\oplus\mathbb{Z}[H]$. The following result of Miyaoka-Nakayama [43, 45] determines the form of the nef and pseudoeffective cones of $\mathbb{P}_{B}(\mathcal{E})$.
@@
    The divisor $kH-\pi^{*}D$ is pseudo-effective (i.e. it is a limit of effective divisors) if and only if $k\geq 0$ and $\deg D\leq k\mu_{1}$.
@@
Although the theorem is stated for divisors, it can be readily extended to $\mathbb{Q}$-divisors, and it is in this form that we will use it.
@@
From Theorem 3.2 (or from the second definition of nefness in Definition 2.4) we see that $\mathcal{E}$ is nef if and only if the smallest slope $\mu_{l}$ is greater or equal to 00. In this case, setting $\mu_{l+1}:=0$, we can reformulate equation (3.1) as follows (see [48, Rem. 3.15]):
$\deg(\mathcal{E})=\sum_{i=1}^{l}r_{i}(\mu_{i}-\mu_{i+1}).$
3.2 The Harder-Narashiman sequence of the Hodge bundle
Let us consider the case of a fibred surface $f\colon S\to B$, and consider its Hodge bundle $\mathcal{E}=f_{*}\omega_{f}$. As $\mathcal{E}$ is nef by Theorem 2.5, we have that $\mu_{l}$ is greater or equal to 00.
Remark 3.5.
Observe that the unitary flat summand $\mathcal{U}$ in (2.7) is the biggest subbundle of $\mathcal{E}$ with degree 00. So, in particular we have that $\mu_{l}=0$ if and only if $\mathcal{U}\not=0$ if and only if $u_{f}\not=0$.
Moreover, if $\mu_{l}=0$, then $\mathcal{E}_{l-1}$ is precisely $\mathcal{A}$.
@@ 3.3 Xiao’s set-up
We follow the set-up given by Xiao in its seminal paper [51]. Let us consider the Hodge bundle $\mathcal{E}=f_{*}\omega_{f}$, and its Harder-Narasimhan sequence as in subsection 3.2. For any $i=1,\ldots l$, we have that the sheaf homomorphism
@@
induces a rational map $\phi_{i}\colon S\dasharrow\mathbb{P}_{B}(\mathcal{E}_{i})$ (see [32, Pro. 7.12]).
Let $H_{i}$ be the tautological divisor on $\mathbb{P}_{B}(\mathcal{E}_{i})$, let $M_{i}:=\phi_{i}^{*}(H_{i})$ on $S$.
@@
Hence, $\phi_{i}^{*}(H_{i}-\mu_{i}\Sigma_{i})=M_{i}-\mu_{i}F$ is a nef $\mathbb{Q}$-divisor on $S$ and $\phi_{i}^{*}(H_{i}-\mu_{1}\Sigma_{i})=M_{i}-\mu_{1}F$ is a pseudoeffective $\mathbb{Q}$-divisor on $S$.
Moreover, for any $i=1,\ldots l$, call $Z_{i}$ the divisorial base locus of $\phi_{i}$, i.e. the effective divisor in $S$ such that $\mathcal{E}_{i}\subseteq f_{*}\omega_{f}(-Z_{i})$ and such that the evaluation homomorphism
@@
is generically surjective. Of course
@@
With the above notations, for any $i=1,\ldots l$, the divisor $K_{f}$ is numerically equivalent to $M_{i}+Z_{i}$,
@@
Let $\tilde{f}=f\circ\rho\colon\widetilde{S}\to B$ be the associated fibration; note that it is non relatively minimal unless $\phi_{i}$ already is a morphism. Let $E_{i}$ be the exceptional divisor of $\rho$.
The morphism $\tilde{\phi}_{i}$ is associated to the line bundle $\mathcal{L}=\rho^{*}\omega_{f}(-Z_{i})\otimes\mathcal{O}_{\widetilde{S}}(-E_{i})$ on $\widetilde{S}$, and we have the surjective sheaf homomorphism ([32, Ch II, Prop 7.11]):
@@
So, from [32, Prop 7.12] we have that $\mathcal{L}\cong\widetilde{\phi_{i}}^{*}(\mathcal{O}_{\mathbb{P}_{B}(\mathcal{E}_{i})}(H_{i}))$. Now, given any irreducible curve $C\subset S$, we have by using the push-pull formula that
$\begin{split}(M_{i}\cdot C)&=(\phi_{i}^{*}(H_{i})\cdot C)=(\tilde{\phi}_{i}^{*}(H_{i})\cdot\rho^{*}(C))=(\rho^{*}(K_{f}-Z_{i})\cdot\rho^{*}(C))+(E_{i}\cdot\rho^{*}(C))=\\ &=(\rho^{*}(K_{f}-Z_{i})\cdot\rho^{*}(C))=(K_{f}-Z_{i}\cdot C),\\ \end{split}$
@@
[Xiao] With the above notations, we have that for any $i=1,\ldots l$:
@@
Note that the divisors $Z_{i}$’s can have vertical components. In case they have fibres as components, say $Z_{i}=bF+Z^{\prime}_{i}$ with $b\geq 0$ and the support of $Z_{i}^{\prime}$ containing no fibre, then we can observe the following:
  • (i)’
@@
  • (ii)’
    $K_{f}-Z^{\prime}_{1}-\mu_{1}F$ is pseudoeffective.
Indeed, for (i)’, let $C$ be any irreducible curve in $S$. If $C$ is vertical then $C\cdot F=0$ and so
@@
For (ii)’ just observe that $K_{f}-Z^{\prime}_{1}-\mu_{1}F=K_{f}-Z_{1}-\mu_{1}F+bF$, so it is certainly pseudoeffective. For this reason in the following we can assume that the $Z_{i}$’s can only have vertical components that are not fibres. We can not suppose that $Z_{i}$ has no vertical components, because (i)’ is not guaranteed if we eliminate those components.
Another approach to get rid of fibres in the base divisors -the one used by Xiao in [51]- is to tensor $\mathcal{E}_{i}$ with a sufficiently ample line bundle $\mathcal{B}$ on $B$ and to consider $L$ the linear subsystem of $|\omega_{f}\otimes f^{*}\mathcal{B}|$ corresponding to sections in $H^{0}(B,\mathcal{E}_{i}\otimes\mathcal{B})$. Then one defines $Z_{i}$ to be the fixed part of $L$ and verifies that this construction is independent on the choice of the ample line bundle $\mathcal{B}$.
Notation 3.11.
Let us call $d_{i}=M_{i}\cdot F$, the degree of the linear series $|V_{i}|$ induced by $M_{i}$ on the general fibre, and let $a_{i}=(Z_{i}\cdot F)=2g-2-d_{i}$. The divisor ${Z_{i}}_{|F}$ on $F$ is the base locus of the linear system $|V_{i}|$ seen as a subcanonical system.
Also call $r_{i}:=\mbox{{rank}}\mathcal{E}_{i}$, so that $|V_{i}|$ is associated to a base point free subcanonical $g^{d_{i}}_{r_{i}-1}$ on $F$.
@@
Observe that $d_{1}\leq d_{2}\leq\ldots d_{l}=2g-2$, and also note that $d_{i}>0$ (i.e. $a_{i}<2g-2$) if $i>1$, because if $i>1$ the series $|V_{i}|$ has dimension $r_{i}-1>0$. Moreover, $d_{1}=0$ if and only if $r_{1}=1$.
3.4 Konno’s idea: a new nef divisor on $S$
Till now we just introduced the framework of Xiao’s technique. Now we turn to an original idea of Konno, which consists in relating the integers $a_{i}$ with the multiplicity of the irreducible components of $Z_{i}$, and then in using the Hurwitz formula to produce new nef divisors on $S$.
@@ 3.4 Konno’s idea: a new nef divisor on $S$
For the reader’s convenience we rielaborate here Konno’s proof of [34, Lemma 2.1]. As in the statement we call $Z_{i}=\sum_{j}m_{j}G_{j}$ be the irreducible decomposition of $Z_{i}$, and let $\alpha_{i}:=\max_{j}\{m_{j}\mid F\cdot G_{j}>0\}$; so, $\alpha_{i}$ is the maximal multiplicity of the components of $Z_{i}$ that are not vertical (note that we already know from Remark 3.10 that we can assume that $Z_{i}$ contains no fibres, but it can contain some component of fibres of $f$). We know from (1) of Theorem 2.2 that $K_{f}$ and from Lemma 3.8 and Proposition 3.9 that $M_{i}-\mu_{i}F$ are nef. Let us now consider any irreducible curve $C\subset S$. If $C$ is not a component of $Z_{i}$, then $C\cdot Z_{i}\geq 0$ and so $(\alpha_{i}K_{f}+M_{i}-\mu_{i}F+Z_{i})\cdot C\geq 0$. Let us now assume that $C=G_{j}$ for some $j$. If $C$ is vertical, $C\cdot F=0$ and so
@@
by the Riemann-Hurwitz formula, where $r(\pi)$ denotes the ramification divisor of $\pi$ on $B$. Now we can rearrange the divisor $\alpha_{i}K_{f}+Z_{i}$ as follows:
@@
$(\alpha_{i}K_{f}+M_{i}-\mu_{i}F+Z_{i})\cdot C=(\alpha_{i}-m_{j})K_{f}\cdot C+m_{j}(K_{f}+C)\cdot C+(Z_{i}-m_{j}C)\cdot C+(M_{i}-\mu_{i}F)\cdot C,$
and one readily verifies that all the contributions are greater or equal to zero, by making the following observations:
@@
The proof is thus concluded. ∎
@@
It is important to remark that this idea of Konno introduces a real novelty with respect to Xiao’s method; indeed, the latter method constructs nef divisors on $S$ by pulling back (extremal) nef divisors on $\mathbb{P}_{B}(\mathcal{E}_{i})$, so it obtains divisors of the form $M_{i}-\mu_{i}F\equiv K_{f}-Z_{i}-\mu_{i}F$. In particular it does not give informations on the nefness of divisors of the form $K_{f}-bF$, unless $Z_{i}=0$.
@@
Notice that $\alpha_{i}\leq a_{i}$, and if $\alpha_{i}=a_{i}$, then necessarily $Z_{i}=a_{i}\Gamma$, where $\Gamma$ is a section of $f$, and moreover $a_{i}\Gamma$ is contained in all the $Z_{j}$ with $j\leq i$.
@@
With the above notations, we have that for any $i=1,\ldots l$,
@@
We consider the rational map $\phi_{i}\colon S\dasharrow\mathbb{P}_{B}(\mathcal{E}_{i})$. As observed above, $K_{f}$ is numerically equivalent to $M_{i}+Z_{i}$, with $Z_{i}$ effective, and $M_{i}-\mu_{i}F$ is nef. By Proposition 3.13, as of course $a_{i}\geq\alpha_{i}$, we have that $(\alpha_{i}+1)K_{f}-\mu_{i}F\equiv\alpha_{i}K_{f}+M_{i}+Z_{i}-\mu_{i}F$ is a nef divisor. So, $((\alpha_{i}+1)K_{f}-\mu_{i}F)\cdot Z_{1}\geq 0$, and ultimately we get that $K_{f}\cdot Z_{i}\geq a_{i}\mu_{i}/(\alpha_{i}+1)$. ∎
We now follow [34, Lemma 2.7] and make explicit the main inequality proved in that argument:
@@
Let $f\colon S\to B$ a relatively minimal non-locally trivial fibred surface of genus $g\geq 2$. Let $a:=a_{1}$, and $\alpha:=\alpha_{1}$. The following inequality holds:
@@
We just consider the first rational map $\phi_{1}\colon S\dasharrow\mathbb{P}_{B}(\mathcal{E}_{1})$. By Lemma 3.16 we have that $K_{f}\cdot Z_{1}\geq a\mu_{i}/(\alpha+1)$.
Now we use a reasoning very close to the classical Xiao’s trick to bound the self-intersection of $K_{f}$:
$\begin{split}K_{f}^{2}&=(K_{f}\cdot M_{1}+Z_{1})=(K_{f}\cdot M_{1}+K_{f}\cdot Z_{1})=(K_{f}\cdot M_{1}-\mu_{1}F)+(K_{f}\cdot\mu_{1}F)+(K_{f}\cdot Z_{1})=\\ &=(K_{f}-\mu_{1}F\cdot M_{1}-\mu_{1}F)+(\mu_{1}F\cdot M-\mu_{1}F)+(K_{f}\cdot\mu_{1}F)+(K_{f}\cdot Z_{1}).\end{split}$ (3.20)
Now, we observe that $K_{f}-\mu_{1}F\equiv M_{1}-\mu_{1}F+Z_{1}$ is an effective divisor, because $M_{1}-\mu_{1}F$ is effective by Theorem 3.2 (as $\mu_{1}\geq\mu_{l}$), and $Z_{1}$ is effective by its very definition. So, the first term $(K_{f}-\mu_{1}F\cdot M_{1}-\mu_{1}F)$ is greater or equal to 00, and we obtain
$\begin{split}K_{f}^{2}\geq\mu_{1}(F\cdot M_{1})+\mu_{1}(K_{f}\cdot F)+(K_{f}\cdot Z_{1})\geq\\ \geq\mu_{1}\left((2g-2-a+2g-2)+\frac{a}{\alpha+1}\right)=\\ =\mu_{1}\left(\frac{(\alpha+1)(4g-4-a)+a}{\alpha+1}\right),\end{split}$ (3.21)
@@
In some cases also these inequalities can be useful, as we will see for instance in the last part of next section, in particular in Theorem 4.30.

### v2

@@ 2.2 Inequalities between the invariants: Xiao’s type of bounds
For arbitrary $b\geq 0$ he proved the bound
@@
He also conjectured that the bound (2.18) holds for any non-trivial fibration. Serrano [37] proved that if $f$ is isotrivial (but not trivial), then (2.18) holds. Cai in [8] proved that if $f$ is non-isotrivial and the general fibre is either hyperelliptic or bielliptic, the same bound holds. In [21], Konno improved the bound (2.19) to:
$q_{f}\leq g\frac{5g-2}{6g-3}.$ (2.20)
Pirola gave a genus $4$ counterexample to Xiao’s conjecture in [34]. More counterexamples have been found by Albano and Pirola in [1]. In [3] Barja, González and Naranjo proved the following: if $f$ is non-isotrivial, then $q_{f}\leq g-c_{f}$. They made the following modified conjecture:
@@
Favale-Naranjo-Pirola [14] proved the stronger inequality $q_{f}\leq g-c_{f}-1$ for families of plane curves of degree $d\geq 5$. It is natural to ask whether these bounds can be extended to the unitary rank $u_{f}$. In the paper [19] the author with González and Torelli proved that : $(i)$ $u_{f}\leq g-c_{f};$ $(ii)$ if the general fibre is a plane curve of degree $d\geq 5$, then $u_{f}\leq g-c_{f}-1.$
@@
If $\mathcal{U}$ has finite monodromy, then up to an ètale base change it becomes trivial, and so coincides with $q_{f}$. However, the unitary summand $\mathcal{U}$ of the Hodge bundle can have infinite monodromy as proved by Catanese and Dettweiler in [9]. Moreover, the modified Xiao’s conjecture does not hold for $u_{f}$: this follows from [9, Remark 38] and from the construction of Lu [29]. See also [19, Conjecture 1.1].
@@ 3 Main technical results
3.1 The Harder-Narasimhan filtration of a vector bundle
The slope of a vector bundle $\mathcal{E}$ over a curve $B$ is defined as $\mu(\mathcal{E}):=\deg(\mathcal{E})/\mbox{{rank}}(\mathcal{E})$. The vector bundle $\mathcal{E}$ is called semistable (resp. stable) if for any vector sub-bundle $\mathcal{F}\subsetneq\mathcal{E}$ we have $\mu(\mathcal{F})\leq\mu(\mathcal{E})$ (resp. $\mu(\mathcal{F})<\mu(\mathcal{E})$).
Let us briefly recall the notion of the Harder-Narasimhan filtration of a vector bundle $\mathcal{E}$ over a curve $B$ [20]. It is the unique filtration of subbundles
$0=\mathcal{E}_{0}\subsetneq\mathcal{E}_{1}\subsetneq\ldots\subsetneq\mathcal{E}_{l}=\mathcal{E}$
@@ 3.1 The Harder-Narasimhan filtration of a vector bundle
  • $\mathcal{E}_{i}/\mathcal{E}_{i-1}$ is a $\mu$-semistable vector bundle for any $i=0,\ldots l$,
  • if we set $\mu_{i}:=\mu(\mathcal{E}_{i}/\mathcal{E}_{i-1})$ for any $i=1,\ldots l$ we have that $\mu_{i}<\mu_{i-1}$.
Note that $\mu_{1}>\mu(\mathcal{E})>\mu_{l}$, unless $\mathcal{E}$ is $\mu$-semistable, in which case $1=l$ and $\mu(\mathcal{E})=\mu_{1}$.
Let us set $r_{i}:=\mbox{{rank}}(\mathcal{E}_{i})$. Setting $r_{0}:=0$, we can express the degree of $\mathcal{E}$ as a combinations of the $\mu_{i}$’s and the $r_{i}$’s:
@@
Let us call $H$ a tautological divisor and $\Sigma$ a fibre of the projective bundle $\pi\colon\mathbb{P}_{B}(\mathcal{E})\to B$. Recall that $\mbox{{Pic}}(\mathbb{P}_{B}(\mathcal{E}))=\pi^{*}\mbox{{Pic}}(B)\oplus\mathbb{Z}[H]$. The following result of Miyaoka-Nakayama [31, 33] determines the form of the nef and pseudo-effective cones of $\mathbb{P}_{B}(\mathcal{E})$.
@@
    The divisor $kH-\pi^{*}D$ is pseudo-effective (i.e., it is a limit of effective divisors) if and only if $k\geq 0$ and $\deg D\leq k\mu_{1}$.
@@
Although the theorem is stated for divisors, it can be readily extended to $\mathbb{Q}$-divisors, and it is this version that we will use.
3.2 The Harder-Narasimhan filtration of the Hodge bundle
Let us consider the case of a fibred surface $f\colon S\to B$, and consider its Hodge bundle $\mathcal{E}=f_{*}\omega_{f}$. As $\mathcal{E}$ is nef by Theorem 2.5, we have that $\mu_{l}$ is greater than or equal to 00.
@@ 3.2 The Harder-Narasimhan filtration of the Hodge bundle
The unitary flat summand $\mathcal{U}$ in (2.7) is the biggest degree 00 subbundle of $\mathcal{E}$. So, in particular we have that $\mu_{l}=0$ if and only if $\mathcal{U}\not=0$ if and only if $u_{f}\not=0$. Moreover, if $\mu_{l}=0$, then $\mathcal{E}_{l-1}$ is precisely $\mathcal{A}$, if $\mu_{l}>0$, then $\mathcal{A}$ is the whole Hodge bundle.
@@ 3.3 Xiao’s set-up
We follow the set-up given by Xiao in his seminal paper [41]. Let us consider the Hodge bundle $\mathcal{E}=f_{*}\omega_{f}$, and its Harder-Narasimhan filtration as in Section 3.2. For any $i=1,\ldots l$, the sheaf homomorphism
@@
induces a rational map $\phi_{i}\colon S\dasharrow\mathbb{P}_{B}(\mathcal{E}_{i})$. Let $H_{i}$ be the tautological divisor on $\mathbb{P}_{B}(\mathcal{E}_{i})$ and let $M_{i}:=\phi_{i}^{*}(H_{i})$ on $S$.
@@
Hence, $\phi_{i}^{*}(H_{i}-\mu_{i}\Sigma_{i})=M_{i}-\mu_{i}F$ is a nef $\mathbb{Q}$-divisor on $S$ and $\phi_{i}^{*}(H_{i}-\mu_{1}\Sigma_{i})=M_{i}-\mu_{1}F$ is a pseudoeffective $\mathbb{Q}$-divisor on $S$. Moreover, for any $i=1,\ldots l$, call $Z_{i}$ the divisorial base locus of $\phi_{i}$, i.e., the effective divisor in $S$ such that $\mathcal{E}_{i}\subseteq f_{*}\omega_{f}(-Z_{i})$ and such that the evaluation homomorphism
@@
is surjective in codimension $2$. Of course
@@
With the above notations, for any $i=1,\ldots l$ the divisor $K_{f}$ is numerically equivalent to $M_{i}+Z_{i}$,
@@
Let $\tilde{f}=f\circ\rho\colon\widetilde{S}\to B$ be the associated fibration; note that it is not relatively minimal unless $\phi_{i}$ already is a morphism. Let $E_{i}$ be the exceptional divisor of $\rho$.
The morphism $\tilde{\phi}_{i}$ is associated to the line bundle $\mathcal{L}=\rho^{*}\omega_{f}(-Z_{i})\otimes\mathcal{O}_{\widetilde{S}}(-E_{i})$ on $\widetilde{S}$, and we have the surjective sheaf homomorphism:
@@
So we have that $\mathcal{L}\cong\widetilde{\phi_{i}}^{*}(\mathcal{O}_{\mathbb{P}_{B}(\mathcal{E}_{i})}(H_{i}))$. Now, given any irreducible curve $C\subset S$, we have by using the push-pull formula that
$\begin{split}M_{i}\cdot C&=\phi_{i}^{*}(H_{i})\cdot C=\tilde{\phi}_{i}^{*}(H_{i})\cdot\rho^{*}(C)=\rho^{*}(K_{f}-Z_{i})\cdot\rho^{*}(C)+E_{i}\cdot\rho^{*}(C)=\\ &=\rho^{*}(K_{f}-Z_{i})\cdot\rho^{*}(C)=(K_{f}-Z_{i})\cdot C,\\ \end{split}$
@@
[Xiao] With the above notations, we have that for any $i=1,\ldots l$
@@
Note that the divisors $Z_{i}$’s can have vertical components. In the case when they have fibres as components, say $Z_{i}=bF+Z^{\prime}_{i}$ with $b\geq 0$ and the support of $Z_{i}^{\prime}$ containing no fibre, then we can observe the following:
  • $(i^{\prime})$
@@
  • ($ii^{\prime}$)
    $K_{f}-Z^{\prime}_{i}-\mu_{1}F$ is pseudoeffective.
Indeed, for $(i^{\prime})$, let $C$ be any irreducible curve in $S$. If $C$ is vertical then $C\cdot F=0$ and so
@@
For $(ii^{\prime})$ just observe that $K_{f}-Z^{\prime}_{i}-\mu_{1}F=K_{f}-Z_{i}-\mu_{1}F+bF$, so it is certainly pseudoeffective. For this reason in the following we can assume that the $Z_{i}$’s can only have vertical components that are not fibres. We cannot suppose that $Z_{i}$ has no vertical components at all, because $(i^{\prime})$ is not guaranteed if we eliminate those components.
Another approach to get rid of fibres in the base divisors –the one used by Xiao in [41]– is to tensor $\mathcal{E}_{i}$ with a sufficiently ample line bundle $\mathcal{B}$ on $B$ and to consider $L$ the linear subsystem of $|\omega_{f}\otimes f^{*}\mathcal{B}|$ corresponding to sections in $H^{0}(B,\mathcal{E}_{i}\otimes\mathcal{B})$. Then one defines $Z_{i}$ to be the fixed part of $L$ and verifies that this construction is independent on the choice of the ample line bundle $\mathcal{B}$.
Notation 3.10.
Let us call $d_{i}=M_{i}\cdot F$, the degree of the linear series $|V_{i}|$ induced by $M_{i}$ on the general fibre, and let $a_{i}=(Z_{i}\cdot F)=2g-2-d_{i}$. The divisor ${Z_{i}}_{|F}$ on $F$ is the base locus of the linear series $|V_{i}|$ seen as a subcanonical system.
Recall that $r_{i}:=\mbox{{rank}}\mathcal{E}_{i}$, for $i=1,\ldots,l$. The linear series $|V_{i}|$ is associated to a base-point-free subcanonical $g^{d_{i}}_{r_{i}-1}$ on $F$.
@@
Observe that $d_{1}\leq d_{2}\leq\ldots\leq d_{l}=2g-2$. Also note that $d_{i}>0$ (i.e., $a_{i}<2g-2$) if $i>1$; indeed, if $i>1$ the linear series $|V_{i}|$ has dimension $r_{i}-1>0$. Moreover, $d_{1}=0$ if and only if $r_{1}=1$.
3.4 A new nef divisor on $S$
Till now we have been working in the framework of Xiao’s technique. Now we turn to an original idea of Konno.
@@ 3.4 A new nef divisor on $S$
For the readers’ convenience we elaborate here Konno’s proof of [21, Lemma 2.1]. Fix $i\in\{1,\ldots,l\}$. Let $Z_{i}=\sum_{j}m_{j}G_{j}$ be the irreducible decomposition of $Z_{i}$, and let
$\alpha_{i}:=\max_{j}\{m_{j}\mid F\cdot G_{j}>0\}.$
In other words, $\alpha_{i}$ is the maximal multiplicity of the non-vertical components of $Z_{i}$. From Remark 3.9 we can assume that $Z_{i}$ contains no fibres, but it can contain some components of fibres. From Theorem 2.2, Lemma 3.7 and Proposition 3.8 we know that $K_{f}$ and $M_{i}-\mu_{i}F$ are nef. Let us now consider any irreducible curve $C\subset S$. If $C$ is not a component of $Z_{i}$, then $C\cdot Z_{i}\geq 0$ and so $(\alpha_{i}K_{f}+M_{i}-\mu_{i}F+Z_{i})\cdot C\geq 0$. Let us now assume that $C=G_{j}$ for some $j$. If $C$ is vertical, $C\cdot F=0$ and so
@@
by the Riemann-Hurwitz formula, where $r(\pi)$ denotes the ramification divisor of $\pi$ on $C$. Now we can rearrange the divisor $\alpha_{i}K_{f}+Z_{i}$ as follows:
@@
$(\alpha_{i}K_{f}+M_{i}-\mu_{i}F+Z_{i})\cdot C=(\alpha_{i}-m_{j})K_{f}\cdot C+m_{j}(K_{f}+C)\cdot C+(Z_{i}-m_{j}C)\cdot C+(M_{i}-\mu_{i}F)\cdot C.$
All the contributions are greater than or equal to zero, by the following observations:
@@
The proposition is thus concluded. ∎
@@
It is important to remark that this idea of Konno introduces a real novelty with respect to Xiao’s method; indeed, the latter method constructs nef divisors on $S$ by pulling back (extremal) nef divisors on $\mathbb{P}_{B}(\mathcal{E}_{i})$, so it obtains divisors of the form $M_{i}-\mu_{i}F\equiv K_{f}-Z_{i}-\mu_{i}F$. In particular it does not give information on the nefness of divisors of the form $K_{f}-bF$, unless $Z_{i}=0$.
@@
Notice that $\alpha_{i}\leq a_{i}$, and if $\alpha_{i}=a_{i}$, then necessarily $Z_{i}=a_{i}\Gamma+Z^{\prime}$, where $\Gamma$ is a section of $f$ and $Z^{\prime}$ is an effective vertical divisor, and moreover $a_{i}\Gamma$ is contained in all the $Z_{j}$ with $j\leq i$.
@@
With the above notations, we have that for any $i=1,\ldots l$
@@
We consider the rational map $\phi_{i}\colon S\dasharrow\mathbb{P}_{B}(\mathcal{E}_{i})$ defined at the beginning of Section 3.3. As proved above, $K_{f}$ is numerically equivalent to $M_{i}+Z_{i}$, with $Z_{i}$ effective and $M_{i}-\mu_{i}F$ nef. By Proposition 3.12 we have that $(a_{i}+1)K_{f}-\mu_{i}F\equiv a_{i}K_{f}+M_{i}+Z_{i}-\mu_{i}F$ is a nef divisor. As $a_{i}\geq\alpha_{i}$, also $(\alpha_{i}+1)K_{f}-\mu_{i}F\equiv\alpha_{i}K_{f}+M_{i}+Z_{i}-\mu_{i}F$ is nef. We thus have $((\alpha_{i}+1)K_{f}-\mu_{i}F)\cdot Z_{i}\geq 0$, and ultimately $K_{f}\cdot Z_{i}\geq a_{i}\mu_{i}/(\alpha_{i}+1)$. ∎
We now derive a general slope inequality:
@@
Let $f\colon S\to B$ be a relatively minimal locally non-trivial fibred surface of genus $g\geq 2$. Let $a:=a_{1}$, and $\alpha:=\alpha_{1}$. The following inequality holds:
@@
We just consider the first rational map $\phi_{1}\colon S\dasharrow\mathbb{P}_{B}(\mathcal{E}_{1})$. By Lemma 3.15 we have that $K_{f}\cdot Z_{1}\geq a\mu_{1}/(\alpha+1)$. Now observe that:
$\begin{split}K_{f}^{2}&=K_{f}\cdot(M_{1}+Z_{1})=K_{f}\cdot M_{1}+K_{f}\cdot Z_{1}=K_{f}\cdot(M_{1}-\mu_{1}F)+K_{f}\cdot\mu_{1}F+K_{f}\cdot Z_{1}=\\ &=K_{f}-\mu_{1}F\cdot(M_{1}-\mu_{1}F)+\mu_{1}F\cdot(M-\mu_{1}F)+(K_{f}\cdot\mu_{1}F+K_{f})\cdot Z_{1}.\end{split}$ (3.19)
Now, we observe that $K_{f}-\mu_{1}F\equiv M_{1}-\mu_{1}F+Z_{1}$ is a pseudo-effective divisor, because $M_{1}-\mu_{1}F$ is pseudo-effective by Theorem 3.2 (as $\mu_{1}\geq\mu_{l}$), and $Z_{1}$ is effective by its very definition. So, the first term $(K_{f}-\mu_{1}F)\cdot(M_{1}-\mu_{1}F)$ is greater than or equal to 00, and we obtain
$\begin{split}K_{f}^{2}&\geq\mu_{1}(F\cdot M_{1})+\mu_{1}(K_{f}\cdot F)+(K_{f}\cdot Z_{1})\geq\quad\quad\quad\quad\quad\\ &\geq\mu_{1}\left((2g-2-a+2g-2)+\frac{a}{\alpha+1}\right)=\mu_{1}\left(\frac{(\alpha+1)(4g-4-a)+a}{\alpha+1}\right),\end{split}$ (3.20)